OpenStudy (anonymous):

1.) A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 �C/m? Show all of the work needed to solve this problem. 2.) The freezing point of a solution will ____________ and the boiling point will ____________ with _____________ molality of a nonvolatile solute. increase, increase, increasing increase, increase, decreasing decrease, increase, increasing decrease, decrease, decreasing

OpenStudy (abmon98):

Number of mole=Mass(g)/Molar Mass(g/mol) 25.5/(12*6)+(1*12)+(16*6)=0.1416667 moles 1 grams represent 1 cm^3 398 grams represent 398 cm^3 Number of Moles=Concentration(mol/dm^3)*Volume(dm^3) Change 398 unit from cm^3 to dm^3 by dividing by 1000 0.141667/0.398=Concentration ΔT = -Kfm ΔT = Change in temperature in °C Kf = freezing point depression constant or cryoscopic constant in °C kg/mol m = molarity of the solute in mol solute/kg solvent.